Accounting 311 In class #1 Based on Case 5-64 in text 5-64. Cost Estimation, CVP Analysis, and Decision Making (LO 5-4, 5, 8)Roy O. Andre, the product manager for Bubbs, is concerned about whether the product will be dropped by the company and has emTable 3 Month Cases 1... 207,000 217,200 214,800 228,000 224,400 237,000 220,200 247,200 238,800 252,600 250,200 259,200 ProdHistogram of Residuals ROC Curve 10000.0 5000.0 + Frequency 0 0.0 -5000.0 - 10000.0 5000 1000 1150000 1250000 - 15000-10000 -Answer the following questions. A regression analysis of the Bubbs production cost data from Table 3 is shown in Table 4. You
Answer #1

1. Number of cases sold = Total Revenue / Selling price per case
= $14682150/5.25 = 2,796,600 cases

2. Variable overhead cost per sales dollar = ($5337500-4221000)/(106750000-76200000) = $0.03655 per sales dollar

3. Variable overhead cost per case = (Cost at highest activity - Cost at lowest activity) / (Highest activity - Lowest Activity)
= ($1272451-1139828)/(259200-207000) = $2.54067 per case

4. Fixed Overhead Cost = $1139828 - 207000 x $2.54067 = $613909

5. Variable overhead cost per case = $2.236 per case

6. Fixed overhead cost nearest to $100 = $682300

As per HOMEWORKLIB RULES we are supposed to answer 4 parts, i have solved 6, so kindly post other questions separately